The measure of one angle of a triangle is 5 less than twice the second angle and the third angle is 45 more than one-half the second angle. find the measure of each angle.
Answer by rfer(10417)

2x-5=75 degrees
x=40 degrees
.5x+45=65 degrees
2x-5+x+.5x+45=180
3.5x=140
x=40